Hi there, I need some help again. Can someone help me solve this? I don't even understand the problem. Thanks! The perimeter of a rectangle is 108m. The length is 9m more then twice the width find the dimensions. What is the length.
Answer by nerdybill(7384)

Hi there, I need some help again. Can someone help me solve this? I don't even understand the problem. Thanks! The perimeter of a rectangle is 108m. The length is 9m more then twice the width find the dimensions. What is the length.
.
Let w = width of rectangle
then
2w+9 = length of rectangle
.
2(width + length) = perimeter
.
2(w + 2w+9) = 108
2(3w+9) = 108
6w+18 = 108
6w = 90
w = 16 m
.
length:
2w+9 = 2(16)+9 = 32+9 = 41 m